) Business and finance. A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?

Answer by stanbon(48551)

A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
--------------
$9 coffee DATA:
Amt of coffee = x lbs ; Value = 9x dollars
---------------------
$12 coffee DATA:
Amt of coffee = "100-x" lbs ; Value = 12(100-x) = 1200-12x dollars
----------------
$11.25 mixture DATA:
Amt of mixture = 100 lbs ; Value = 11.25*100 = 1125 dollars
--------------
EQUATION:
value + value = value of the mixture
9x + 1200-12x = 1125
-3x = -75
x = 25 lbs ( amt of $9 coffee in the mixture)
100-x = 75 ( amt of $12 coffee in the mixture)
===============
Cheers,
Stan H.